OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (cwrw238):

find y-intercept by putting x = 0 and doing the math and for x intercept do same using y = 0

OpenStudy (cwrw238):

i;ll do x-intercept for you y= 0 so 3x - 2(0) = 6 3x = 6 x = 6/3 = 2